OpenStudy (anonymous):

Write this equation in slope-intercept form 6x+4y=0

OpenStudy (anonymous):

I have solved it already not just sure of myself 6x+4y=0 6x-6x=0-6x 4y=0-6x 4y/4=0/4-6x y=0-6x/4 y=-0x/2+2

OpenStudy (tkhunny):

?? Not quite sure what it is you are doing. 6x + 4y = 0 Subtract 6x 4y = -6x Divide by 4 y = (-6x)/4 = -3x/2 That's it. y = (-3/2)x <== You may add the "+0" if you like.

OpenStudy (anonymous):

To the equation above in slope intercept form

OpenStudy (tkhunny):

Does it look like y = mx + b when written as y = (-3/2)x + 0 ??

OpenStudy (anonymous):

Write this equation in slope-intercept form 6x+4y=0